Add: a + b – 3, b – a + 3, a – b + 3 Add: a + b – 3, b – a + 3, a – b + 3

Add: 
 a + b – 3, b – a + 3, a – b + 3
Add: 
 a + b – 3, b – a + 3, a – b + 3

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

When term have the same algebraic factors, they are like terms. Then, we have to add the like terms
= a + b–3 + (b–a + 3) + (a–b + 3)
= a + b–3 + b–a + 3 + a–b +3
= a–a + a+ b + b–b–3 + 3 + 3
= a (1–1 + 1) + b (1 + 1 - 1) + (-3 + 3 + 3)
= a (2 -1) + b (2 -1) + (-3 + 6)
= a (1) + b (1) + (3)
= a + b+ 3

Thanks